OpenStudy (anonymous):

Let x ∈ Z. Prove that if 5x - 7 is odd, then 9x + 2 is even. @satellite73 @ganeshie8 @myininaya @UnkleRhaukus @RadEn @robtobey @chmvijay @Luigi0210 @eliassaab @genius12 @Vincent-Lyon.Fr @Preetha @Kainui @ehuman @wolfe8 @lucaz @INeedHelpPlease? @khadeeja @thadyoung @A_clan @tester97 @Andras @Confusionist @linh412986 @BlackLabel @leozap1 @JMark @Microrobot @zacharyf @Rubio101 @NaomiBell1997 @MeganEdward @theballer225 @link,zoro @Goldenmoon @CrayolaCrayon_ @Cutefriendzoned @coffeeismylife @christopher_miller420 @soumyamohan @zubhanwc3 @Ahmed47 @becca9898 @Marharg10 @kels428 @haunting_beauty @kauf

OpenStudy (luigi0210):

Everyone's sleeping and I feel alone ;-;

ganeshie8 (ganeshie8):

5x-7 is odd => 5x-7 = 1 (mod 2) x = 0 (mod 2) -----------(1)

OpenStudy (vincent-lyon.fr):

Well, is it necessary to tag so many people? Show your work, please.

ganeshie8 (ganeshie8):

9x +2 = x = 0 (mod 2) so 9x +2 is even thats one method

OpenStudy (tester97):

this is your 3rd time tagging me in a mass tag and i frankly do not find this amusing so it would be nice and the right thing to do if you ceased from this Tom Foolery because franlky im ticked off -_-

ganeshie8 (ganeshie8):

so much hatred for the mass tagging addicts :)

OpenStudy (luigi0210):

Can I mass tag next?

ganeshie8 (ganeshie8):

idk, i never felt it necessary to masstag ppl... maybe im here for long time i knw exactly who can help me wid my questions humm

OpenStudy (anonymous):

@ganeshie8 Can you also prove it using a lemma (Pit-stop proof)

ganeshie8 (ganeshie8):

never heard of such proof in number theory; what is it ?

OpenStudy (anonymous):

Try not to prove the second statement directly from the first statement, but rather prove something else from the first statement and use that proof to prove the second statement.

ganeshie8 (ganeshie8):

If you follow the modular proof above, thats exactly what we did. from the first statement we got x = 0 (mod 2). this proves x is even. we have used that in second statement and concluded

OpenStudy (anonymous):

Fair enough. <3

ganeshie8 (ganeshie8):

:) your questions are nice ones... but ppl are getting mad cuz of masstag it seems -_-

OpenStudy (anonymous):

Other people will only benefit from its greatness. 1 Corinthians 2:9.

OpenStudy (tester97):

quoting the bible are we?

ganeshie8 (ganeshie8):

lol lets establish some harmony ;) http://batish.com/catalog/inst/Harmonium/

OpenStudy (anonymous):

No I'm not, the actual verse is much greater: However, as it is written: “What no eye has seen, what no ear has heard, and what no human mind has conceived the things God has prepared for those who love him—"

OpenStudy (ikram002p):

written where ?

OpenStudy (anonymous):

1 Corinthians 2:9

OpenStudy (ikram002p):

see , in Hadeth for muslim there is somthing semi its like this What no eye has seen, what no ear has heard, and what no human mind has conceived the things God has prepared for those who do good things

OpenStudy (anonymous):

@ganeshie8 http://www.chick.com/m/reading/tracts/readtract.asp?stk=0001

ganeshie8 (ganeshie8):

scary lol :o

OpenStudy (anonymous):

@FutureMathProfessor give us a tougher question pls xP

OpenStudy (anonymous):

Ok gimme a little bit

OpenStudy (anonymous):

By the way I'm overdosing on caffeine again like last night. Haven't slept in 42 hours now yolo

OpenStudy (anonymous):

wtf???? ur joking right? @FutureMathProfessor 42 hours? that's actually really bad lol.

OpenStudy (anonymous):

@genius12 When I was at college I was asleep 11-7 for four months nonstop. I Think I can live this horrible lifestyle for a week until I go back :)

OpenStudy (anonymous):

ok gimme harda question nowsss